OpenStudy (anonymous):

Find the smaller root for x^2+7=15

OpenStudy (campbell_st):

subtract 7 from both sides x^2 = 8 take the square root \[x = \pm \sqrt{8}\] or \[x = \pm2\sqrt{2}\]

OpenStudy (anonymous):

the answer is -2 ?

OpenStudy (anonymous):

hello ??

OpenStudy (thomas5267):

Do you know the quadratic formula?

OpenStudy (anonymous):

no ?

OpenStudy (campbell_st):

if the answer is 2... there is a typo in the question... perhaps 2x^2 + 7 = 15

OpenStudy (anonymous):

i put 2 and -2 they are both wrong

OpenStudy (campbell_st):

but the smaller is \[x = -\sqrt{8}\] or \[x = -2\sqrt{?}\] they have the same value

OpenStudy (campbell_st):

oops \[x = -2\sqrt{2}\]